OpenStudy (anonymous):

find the equation, in standard form, of the line passing through the points (2,-3) and (4,2)? I got .4 for b and 2/5 for slope but after that I can't seem to get the right answer. If anyone can help I'd appreciate it.

OpenStudy (anonymous):

\[m = \frac{ y2 - y1 }{ x2 - x1 } = \frac{ 2 - (-3) }{4- 2 } = \frac{ 5 }{ 2 }\]

OpenStudy (anonymous):

\[y = \frac{ 5 }{ 2 }x + b...substituting.gives -3 = \frac{ 5 }{ 2 }(2) + b...solving.for.b.gives...b = -8\]

OpenStudy (anonymous):

Thanks, I finally got it :)